ms.boy's age pus twice arun's age is 85. the difference of a square is 2800. what are the ages?
Answer by ankor@dixie-net.com(22740)

.boy's age plus twice arun's age is 85.
the difference of a square is 2800.
:
let b = boy's age
let a = A's age
:
write an equation for each statement (as I understand it)
:
"boy's age plus twice arun's age is 85."
b + 2a = 85
b = (-2a+85)
:
"the difference of their squares is 2800."
b^2 - a^2 = 2800
replace b with (-2a+85)
(-2a+85)^2 - a^2 = 2800
4a^2 - 340a + 7225 - a^2 = 2800
3a^2 - 340a + 7225 - 2800 = 0
3a^2 - 340a + 4425 = 0
Use the quadratic formula here, but this will factor to
(3a-295)(a-15) = 0
a = 15 is Arun's age (the only reasonable solution)
then
b = -2(15) + 85
b = 55 yrs is boy's age
:
:
Check this on your calc: 55^2 - 15^2